An investor has decided to purchase a fixed interest bond at issue which offers coupons of 5% per annum, payable half-yearly. The bond will be redeemed at 110% in 10 years’ time.
The investor pays income tax at 30% and capital gains tax at 20%.

Calculate the price per £100 nominal if the investor is to obtain a yield of 6% per annum. (Total 6 marks)
1
Expert's answer
2020-04-06T17:49:08-0400

сalculate according to the formula:

"P=\frac{C}{r}(1-\frac{1}{(1+r)^{mn}})+\frac{N}{(1+r)^{mn}}=\frac{5}{0.06}(1-\frac{1}{(1+0.06)^{20}})+\frac{105}{(1+0.06)^{20}}=83.33(1-\frac{1}{(1.06)^{20}})+\frac{105}{(1.06)^{20}}=57.34+32.74=90.08"


"С=100\times 0.05=5"

"N=100\times1.05=105"

"mn=10\times2=20"